# Approximations for the number of maxima and near-maxima in independent data

### v1

@@ 1 Introduction and main results
In this we are motivated by the work of Brands et al. [3], who show that, while $K_{n}$ does not in general converge to a distributional limit as $n\to\infty$, there are nevertheless good approximations to the distribution of $K_{n}$. Motivated by a coin-tossing problem of Räde [12], Brands et al. pay particular attention to the case where $X$ is geometrically distributed, and observe that in this case if the parameter $p$ is constant (i.e., independent of $n$) the distribution of $K_{n}$ is close to logarithmic, while for other choices of $p$ it may be close to Poisson. Our first principal aim here is to complement these observations with explicit error bounds in total variation distance for the approximation of the distribution of $K_{n}$ by either a logarithmic or Poisson distribution. The number of maxima in the case where $X$ is geometrically distributed has also been paid particular attention by Kirschenhofer and Prodinger [8] and Olofsson [9], and we refer the interested reader to the work of Eisenberg [7] and Bruss and Grübel [5] and references therein for a discussion of related results in a more general discrete setting. We will treat the general case here, but use the geometric distribution as a motivating and illustrative example.
In the setting where $X$ is absolutely continuous, the situation is somewhat different. Here we consider instead the number of observations within some threshold $a$ from the maximum. Pakes and Steutel [11] establish a limiting (shifted) mixed Poisson distribution for this quantity as $n\to\infty$, and note that this mixed Poisson simplifies to a geometric distribution in the case where $X$ is in the maximum domain of attraction of a Gumbel distribution. This geometric limit for observations close to the maximum is generalised to a negative binomial limit for observations close to one of the order statistics of our data by Pakes and Li [10]. Our second principal aim in this note is to provide explicit error bounds (again in total variation distance) in the negative binomial approximation of the number of data points close to the order statistics of our sample in this absolutely continuous setting. We will formulate this problem more precisely later in this section, after first stating and illustrating our main results in the discrete case.
All of our approximations and error bounds will be given in the total variation distance, defined for random variables $K$ and $L$ supported on +=$\mathbb{Z}^{+}=\{0,1,\ldots\}$ by
@@
We use the remainder of this section to present our main results in the discrete setting (Theorems 1 and 3 below) and absolutely continuous case (in Theorem 5), and to further provide examples to illustrate these results. Proofs of these main results are deferred to Sections 24. These proofs make use of Stein’s method, adapted here for use with a logarithmic target distribution for the first time and also used to establish new results in the negative binomial approximation of mixed binomial distributions. We will also make use of well-known results on Poisson approximation.
@@ 1.1 The discrete case
for $k=1,\ldots,n$. From this, expressions for the factorial moments of $K_{n}$ follow easily. In particular, letting $(k)_{\ell}=k(k-1)\cdots(k-\ell+1)$ denote the falling factorial, we have that
@@
The proof of Theorem 1(a) is given in Section 2 below, while part (b) is proved in Section 3. Before moving on to state our Poisson approximation results, we illustrate the upper bounds of Theorem 1 in our motivating example of geometrically distributed data.
@@
Suppose that $\mathbb{P}(X=j)=p(1-p)^{j-1}$ for $j=1,2,\ldots$ and some $p\in(0,1)$, so that $X\sim\text{Geom}(p)$ has a geometric distribution with parameter $p$. In this example we can check numerically that the upper bound of part (a) of Theorem 1 is superior to that of part (b), so we focus the former result here. We nevertheless include part (b) in the statement of Theorem 1, since there may be other examples where it is useful and since its proof uses tools very similar to those needed for Theorems 3 and 5 below.
In the geometric setting we can easily check that the parameter $\alpha$ of our approximating logarithmic distribution is given by $\alpha=p$: using (2) and (3) we have
$1-\alpha=\frac{\sum_{j=1}^{\infty}p(1-p)^{j-1}[1-(1-p)^{j-1}]^{n-1}}{\sum_{j=1}^{\infty}p(1-p)^{j-1}[1-(1-p)^{j}]^{n-1}}=(1-p)\frac{\sum_{j=2}^{\infty}p(1-p)^{j-1}[1-(1-p)^{j-1}]^{n-1}}{\sum_{j=1}^{\infty}p(1-p)^{j}[1-(1-p)^{j}]^{n-1}}=1-p\,.$
Numerical illustration of the upper bound of Theorem 1(a) in this geometric example is given in Figure 1 for the case $n=20$ and various small-to-moderate values of $p$. We focus on such values of $p$ here since the proof of Theorem 1(a) is geared towards this case: see Remark 10 below. Other values of $n$ give upper bounds which are very similar to those shown on Figure 1.
@@
$d_{\text{TV}}(K_{n},Y)\leq\frac{\sqrt{\mathbb{E}[(K_{n})_{2}]-\mathbb{E}[K_{n}](\mathbb{E}[K_{n}]-1)}}{2\mathbb{E}[K_{n}]}+\sqrt{\frac{\mathbb{E}[K_{n}]}{4\mathbb{E}[(K_{n})_{2}]}}+\frac{(n-1)\mathbb{E}[(K_{n})_{3}]}{(n-2)\mathbb{E}[(K_{n})_{2}]}-\frac{(n-2)\mathbb{E}[(K_{n})_{2}]}{(n-1)\mathbb{E}[K_{n}]}\,.$
@@ 1.2 The absolutely continuous case
We turn now to the case where $X$ is a real-valued random variable with probability density function $f(x)$ and distribution function $F(x)=\int_{-\infty}^{x}f(y)\,\text{d}y$ for $x\in\mathbb{R}$. Following Pakes and Li [10], we consider the more general setting of negative binomial approximation for the number of observations within a certain distance of one of the order statistics of our sample $X_{1},\ldots,X_{n}$. To that end, we let $X_{1:n}<\cdots<X_{n:n}$ denote the order statistics of our sample, and for $\ell=1,\ldots,n$ we define
$K_{n}(a,\ell)=|\left\{i\in\{1,\ldots,n\}:X_{n-\ell+1:n}-a<X_{i}<X_{n-\ell+1:n}\right\}|\,,$ (4)
the number of our $n$ observations that are within a distance $a$ of the $\ell$th order statistic. Pakes and Li (see Theorem 1 of [10]) use a mixed binomial representation of $K_{n}(a,\ell)$ to show that if $X$ is in the maximum domain of attraction of a Gumbel law, then $K_{n}(a,\ell)$ may be well approximated by a negative binomial random variable. We aim to complement this result with explicit error bounds in total variation distance.
@@
Let $X$ have a Gumbel distribution with density function $f(x)=e^{-x+e^{-x}}$ and distribution function $F(x)=e^{-e^{-x}}$ for $x\in\mathbb{R}$, which has mean equal to $0.57721\ldots$, the Euler–Mascheroni constant, and standard deviation $\pi/\sqrt{6}$. We consider geometric approximation for the number of $n$ i.i.d. samples from $X$ which are within distance $a$ of their maximum; that is, we take $\ell=1$ in Theorem 5. Note that the maximum of our $n$ data points has itself a Gumbel distribution, shifted $\log(n)$ units to the right. With this choice of $X$ we have that
$M_{1}$ $=n\int_{-\infty}^{\infty}(e^{-e^{-x}})^{n-1}\left(1-\frac{e^{-e^{-(x-a)}}}{e^{-e^{-x}}}\right)e^{-x+e^{-x}}\,\text{d}x=\frac{e^{a}-1}{n+e^{a}-1}\,,\text{ and}$
$M_{2}$ $=n\int_{-\infty}^{\infty}(e^{-e^{-x}})^{n-1}\left(1-\frac{e^{-e^{-(x-a)}}}{e^{-e^{-x}}}\right)^{2}e^{-x+e^{-x}}\,\text{d}x=\frac{2(e^{a}-1)^{2}}{(n+e^{a}-1)(n+2e^{a}-2)}\,.$
Using the mixed binomial representation of $K_{n}(a,1)-1$ (see Lemma 1 of [10] and Section 3.2 below) we have that
@@
$d_{\text{TV}}(K_{n}(a,1)-1,Z)$ $\leq\frac{(n-1)(e^{a}-1)}{n+e^{a}-1}\bigg{(}\frac{(n-1)(e^{a}-1)}{ne^{a}}+$
$\qquad\qquad\qquad\left(1-\frac{(n-1)(e^{a}-1)}{ne^{a}}\right)\left[\frac{2(n-2)(e^{a}-1)}{n+2e^{a}-2}-\frac{(n-3)(e^{a}-1)}{n+e^{a}-1}\right]\bigg{)}$
$=\frac{(n-1)(e^{a}-1)^{2}}{e^{a}(n+e^{a}-1)}\left(1+\frac{n-2}{n+2e^{a}-2}\right)\,.$ (5)
We can see that unfortunately our upper bound is not strong enough to converge to zero as $n\to\infty$ for any fixed $a$, in which setting we know we have a geometric limit thanks to Theorem 1 of [10]. The coupling used in the proof of Theorem 5 does not seem to be strong enough to establish this. Nevertheless, our upper bound does tend to zero if $a=a(n)\to 0$ as $n\to\infty$, and also gives reasonable numerical bounds on $d_{\text{TV}}(K_{n}(a,1)-1,Z)$, even for moderate values of $n$ and $a$. This is illustrated in Figure 2, which gives the upper bound (5) in the cases $n=20$ and $n=100$ for various values of $a$.
@@
Let $X\sim\text{U}(0,b)$ have a uniform distribution on the interval $(0,b)$, so that $f(x)=1/b$ and $F(x)=x/b$ for $0\leq x\leq b$. The uniform distribution is in the maximum domain of attraction of a Weibull (not a Gumbel) distribution, so although we do not expect a negative binomial limit for $K_{n}(a,\ell)-1$ in the case of fixed $b$ here, we may nevertheless apply our theorem to give explicit upper bounds in the approximation of $K_{n}(a,\ell)$ which may be useful both numerically and in determining conditions on $a=a(n)$, $b=b(n)$ and $\ell=\ell(n)$ under which we may see a negative binomial limit as $n\to\infty$. With our choice of $X$,
$M_{1}$ $=\frac{an}{b^{n-\ell+1}}\binom{n-1}{\ell-1}\int_{0}^{b}x^{n-\ell-1}\left(1-\frac{x}{b}\right)^{\ell-1}\,\text{d}x=\frac{an}{b(n-\ell)}\,,\text{and}$
$M_{2}$ $=\frac{a^{2}n}{b^{n-\ell+1}}\binom{n-1}{\ell-1}\int_{0}^{b}x^{n-\ell-2}\left(1-\frac{x}{b}\right)^{\ell-1}\,\text{d}x=\frac{a^{2}n(n-1)}{b^{2}(n-\ell)(n-\ell-1)}\,.$
As in the Gumbel example above, we use the mixed binomial representation of $K_{n}(a,\ell)-1$ to get that
$\mathbb{E}[K_{n}(a,\ell)]-1=(n-\ell)M_{1}=\frac{an}{b}\,,$
so that our approximating negative binomial distribution has $\beta=\frac{an}{an+b\ell}$. With $Z\sim\text{NB}(\ell,1-\beta)$, our Theorem 5 then gives the upper bound
$d_{\text{TV}}(K_{n}(a,\ell)-1,Z)$ $\leq\frac{a}{b\ell}(1-(1-\beta)^{\ell})\left[n+\ell\left(n-1-\frac{n(n-\ell-2)}{n-\ell}\right)\right]$
$=\frac{a}{b\ell}(1-(1-\beta)^{\ell})\left(n+\frac{\ell(n+\ell)}{n-\ell}\right)\,.$
Then, for example, for a fixed $\ell$ we may obtain a negative binomial limit for $K_{n}(a,\ell)-1$ if $a/b\to 0$ fast enough as $n\to\infty$.

### v2

@@ 1 Introduction and main results
Eisenberg [7] discusses applications of this statistic to sporting competitions, modelling the following scenario: assuming a set of players of equal ability, how many are tied for the record performance? Similarly, $K_{n}$ finds applications in the reliability of systems composed of $n$ independent components with identically distributed lifetimes, and in the output of randomised selection algorithms (see [5] for a brief discussion of this final application).
Here we are motivated by the work of Brands et al. [3], who show that, while $K_{n}$ does not in general converge to a distributional limit as $n\to\infty$, there are nevertheless good approximations to the distribution of $K_{n}$. Motivated by a coin-tossing problem of Räde [12], Brands et al. pay particular attention to the case where $X$ is geometrically distributed with parameter $p$, and observe that if $p$ is constant (i.e., independent of $n$) the distribution of $K_{n}$ is close to logarithmic, while for other choices of $p$ it may be close to Poisson. Our first principal aim here is to complement these observations with explicit error bounds in total variation distance for the approximation of the distribution of $K_{n}$ by either a logarithmic or Poisson distribution, and thus obtain simple approximations relevant to the applications discussed above together with quantitative error bounds. The number of maxima in the case where $X$ is geometrically distributed has also been paid particular attention by Kirschenhofer and Prodinger [8] and Olofsson [9], and we refer the interested reader to the work of Eisenberg [7] and Bruss and Grübel [5] and references therein for a discussion of related results in a more general discrete setting. We will treat the general case here, but use the geometric distribution as a motivating and illustrative example.
In the setting where $X$ is absolutely continuous, the situation is somewhat different. Here we consider instead the number of observations within some threshold $a$ from the maximum. Pakes and Steutel [11] establish a limiting (shifted) mixed Poisson distribution for this quantity as $n\to\infty$, and note that this mixed Poisson simplifies to a geometric distribution in the case where $X$ is in the maximum domain of attraction of a Gumbel distribution. This geometric limit for observations close to the maximum is generalised to a negative binomial limit for observations close to one of the order statistics of our data by Pakes and Li [10]. Our second principal aim in this note is to provide explicit error bounds (again in total variation distance) in the negative binomial approximation of the number of data points close to the order statistics of our sample in this absolutely continuous setting, relevant to applications analogous to those of $K_{n}$ above. We will formulate this problem more precisely later in this section, after first stating and illustrating our main results in the discrete case.
All of our approximations and error bounds will be given in the total variation distance, defined for random variables $K$ and $L$ supported on $\mathbb{Z}^{+}=\{0,1,\ldots\}$ by
@@
We use the remainder of this section to present our main results in the discrete setting (Theorems 1 and 3 below) and absolutely continuous case (in Theorem 5), and to further provide examples to illustrate these results. Proofs of these main results are deferred to Sections 24. These proofs make use of Stein’s method, adapted here for use with a logarithmic target distribution for the first time and also used to establish new results in the negative binomial approximation of mixed binomial distributions. We will also make use of well-known results on Poisson approximation. Some concluding remarks are given in Section 5.
@@ 1.1 The discrete case
for $k=1,\ldots,n$.
Although the mass function (2) appears relatively simple, it includes a periodic component that adds complexity to its asymptotics; see Proposition 1 of [8] for asymptotic results in the case where $X$ has a geometric distribution. This result motivates us to derive approximations in total variation distance for $K_{n}$, quantifying the error in simple approximations for expectations of bounded functions of $K_{n}$, not only the mass function and moments as in [8].
From (2), expressions for the factorial moments of $K_{n}$ follow easily. In particular, letting $(k)_{\ell}=k(k-1)\cdots(k-\ell+1)$ denote the falling factorial, we have that
@@
The proof of Theorem 1(a) is given in Section 2 below, while part (b) is proved in Section 3. Note that we typically expect the upper bound of part (a) to be superior to that of part (b); see Remark 13 below. We nevertheless include part (b) in the statement of Theorem 1, since there may be examples where it is useful and since its proof uses tools very similar to those needed for Theorems 3 and 5 below.
Before moving on to state our Poisson approximation results, we illustrate the upper bounds of Theorem 1 in our motivating example of geometrically distributed data.
@@
We let $X\sim\text{Geom}(p)$ have a geometric distribution with parameter $p\in(0,1)$, that is, $\mathbb{P}(X=j)=p(1-p)^{j-1}$ for $j=1,2,\ldots$. In this example we can check numerically that the upper bound of part (a) of Theorem 1 is superior to that of part (b), so we focus on the former result here. In the geometric setting we can easily check that the parameter $\alpha$ of our approximating logarithmic distribution is given by $\alpha=p$: using (2) and (3) we have
$1-\alpha$ $=\frac{\sum_{j=1}^{\infty}p(1-p)^{j-1}[1-(1-p)^{j-1}]^{n-1}}{\sum_{j=1}^{\infty}p(1-p)^{j-1}[1-(1-p)^{j}]^{n-1}}$
$=(1-p)\frac{\sum_{j=2}^{\infty}p(1-p)^{j-1}[1-(1-p)^{j-1}]^{n-1}}{\sum_{j=1}^{\infty}p(1-p)^{j}[1-(1-p)^{j}]^{n-1}}=1-p\,.$
Numerical illustration of the upper bound of Theorem 1(a) in this geometric example is given in Figure 1 for the case of $n=20$ and various small-to-moderate values of $p$. We focus on such values of $p$ here since the proof of Theorem 1(a) is geared towards this case: see Remark 10 below. Other values of $n$ give upper bounds which are very similar to those shown on Figure 1. As expected (see (14) in Remark 10 below), our upper bound is very close to
$2p\left(\frac{1}{1-p}-(1-p)\right)$
for those values of $p$ that we consider here.
We note, however, that even for moderate $p$ there seems to be significant room for improvement in the upper bounds we obtain. For $n=20$ and for several values of $p$ between $0.1$ and $0.2$, we simulated $10^{7}$ independent copies of $K_{n}$ for each $p$ and estimated the total variation distance $d_{\text{TV}}(K_{n},L)$, with $L$ as in Theorem 1(a), corresponding to that choice of $p$. In each case we obtained an estimate of order $10^{-5}$, significantly smaller than the upper bounds shown on Figure 1 for these values of $p$.
@@
$d_{\text{TV}}(K_{n},Y)\leq\frac{\sqrt{\mathbb{E}[(K_{n})_{2}]-\mathbb{E}[K_{n}](\mathbb{E}[K_{n}]-1)}}{2\mathbb{E}[K_{n}]}+\sqrt{\frac{\mathbb{E}[K_{n}]}{4\mathbb{E}[(K_{n})_{2}]}}\\ +\frac{(n-1)\mathbb{E}[(K_{n})_{3}]}{(n-2)\mathbb{E}[(K_{n})_{2}]}-\frac{(n-2)\mathbb{E}[(K_{n})_{2}]}{(n-1)\mathbb{E}[K_{n}]}\,.$
@@ 1.2 The absolutely continuous case
We turn now to the case where $X$ is a real-valued random variable with probability density function $f(x)$ and distribution function $F(x)=\int_{-\infty}^{x}f(y)\,\text{d}y$ for $x\in\mathbb{R}$. Following Pakes and Li [10], we consider the more general setting of negative binomial approximation for the number of observations within a certain distance of one of the order statistics of our sample $X_{1},\ldots,X_{n}$. To that end, we let $X_{1:n}<\cdots<X_{n:n}$ denote the order statistics of our sample, and for $\ell=1,\ldots,n-1$ we define
$K_{n}(a,\ell)=|\left\{i\in\{1,\ldots,n\}:X_{n-\ell+1:n}-a<X_{i}\leq X_{n-\ell+1:n}\right\}|\,$, (4)
[…]
@@
Consider the case where $a$ and $\ell$ are fixed (i.e., independent of $n$). The mixed binomial representation of $K_{n}(a,\ell)-1$ (see Lemma 1 of [10] and Section 3.2 below) gives that
$\mathbb{E}[K_{n}(a,\ell)]-1=(n-\ell)M_{1}\,$,
so that Theorem 5 gives an upper bound of order $O(nM_{1}[1+\frac{nM_{2}}{M_{1}}])=O(nM_{1}+n^{2}M_{2})$. We will see, for example, in Example 6 below that when the $X_{i}$ have a Gumbel distribution, $M_{j}=O(n^{-j})$ for $j=1,2$ in this setting, so that our upper bound is of order $O(1)$. We note, however, that in this example and more generally, if $a$ and $\ell$ are allowed to depend on $n$ the order of the upper bound of Theorem 5 will vary.
@@
Let $X$ have a Gumbel distribution with density function $f(x)=e^{-x-e^{-x}}$ and distribution function $F(x)=e^{-e^{-x}}$ for $x\in\mathbb{R}$, which has mean equal to $0.57721\ldots$, the Euler–Mascheroni constant, and standard deviation $\pi/\sqrt{6}$. We consider a sample of $n$ i.i.d. observations from $X$, and establish a geometric approximation bound for the number of these which are within distance $a$ of the sample maximum; that is, we take $\ell=1$ in Theorem 5. Note that the maximum of our $n$ data points has itself a Gumbel distribution, shifted $\log(n)$ units to the right. With this choice of $X$ we have that
$M_{1}$ $=n\int_{-\infty}^{\infty}(e^{-e^{-x}})^{n-1}\left(1-\frac{e^{-e^{-(x-a)}}}{e^{-e^{-x}}}\right)e^{-x-e^{-x}}\,\text{d}x=\frac{e^{a}-1}{n+e^{a}-1}\,,\text{ and}$
$M_{2}$ $=n\int_{-\infty}^{\infty}(e^{-e^{-x}})^{n-1}\left(1-\frac{e^{-e^{-(x-a)}}}{e^{-e^{-x}}}\right)^{2}e^{-x-e^{-x}}\,\text{d}x=\frac{2(e^{a}-1)^{2}}{(n+e^{a}-1)(n+2e^{a}-2)}\,.$
Again using the mixed binomial representation of $K_{n}(a,1)-1$ we have that
@@
$d_{\text{TV}}(K_{n}(a,1)-1,Z)\leq\frac{(n-1)(e^{a}-1)}{n+e^{a}-1}\bigg(\frac{(n-1)(e^{a}-1)}{ne^{a}}+\\ \left(1-\frac{(n-1)(e^{a}-1)}{ne^{a}}\right)\left[\frac{2(n-2)(e^{a}-1)}{n+2e^{a}-2}-\frac{(n-3)(e^{a}-1)}{n+e^{a}-1}\right]\bigg)\,$,
which simplifies to give
$d_{\text{TV}}(K_{n}(a,1)-1,Z)\leq\frac{(n-1)(e^{a}-1)^{2}}{e^{a}(n+e^{a}-1)}\left(1+\frac{n-2}{n+2e^{a}-2}\right)\,.$ (5)
We can see that unfortunately our upper bound is not strong enough to converge to zero as $n\to\infty$ for any fixed $a$, in which setting we know we have a geometric limit thanks to Theorem 1 of [10]. The coupling used in the proof of Theorem 5 does not seem to be strong enough to establish this. Nevertheless, our upper bound does tend to zero if $a=a(n)\to 0$ as $n\to\infty$, and also gives reasonable numerical bounds on $d_{\text{TV}}(K_{n}(a,1)-1,Z)$, even for moderate values of $n$ and $a$. This is illustrated in Figure 2, which gives the upper bound (5) in the cases $n=20$ and $n=100$ for various values of $a$. The numerical upper bounds of Figure 2 do, however, leave some room for improvement, overestimating the total variation distance by at least an order of magnitude for the examples illustrated. Using $10^{7}$ independent simulations of $K_{n}(a,1)$ for $n=20$ and each $a=0.05,0.1,0.15,0.2$, we estimate $d_{\text{TV}}(K_{n}(a,1)-1,Z)$ as in (5) to be $0.0002,0.0008,0.0014,0.0024$, respectively, for each value of $a$.
@@
Let $X\sim\text{U}(0,1)$ have a uniform distribution on the interval $(0,1)$, so that $f(x)=1$ and $F(x)=x$ for $0\leq x\leq 1$. The uniform distribution is in the maximum domain of attraction of a Weibull (not a Gumbel) distribution, so although we do not expect a negative binomial limit for $K_{n}(a,\ell)-1$ in the case of fixed $a$ here, we may nevertheless apply our theorem to give explicit upper bounds in the approximation of $K_{n}(a,\ell)-1$ which may be useful both numerically and in determining conditions on $a=a(n)$ and $\ell=\ell(n)$ under which we may see a distribution which is close to a negative binomial. We assume that $a=a(n)\in(0,1)$ for each $n$, and for simplicity we will here assume that $\ell$ is fixed (i.e., independent of $n$). With our choice of $X$ we have that
[…]
so that for $j=1,2$ we have
$M_{j}=n\binom{n-1}{\ell-1}\left[\int_{0}^{a}x^{n-\ell}\left(1-x\right)^{\ell-1}\,\text{d}x+a^{j}\int_{a}^{1}x^{n-\ell-j}\left(1-x\right)^{\ell-1}\,\text{d}x\right]\,.$
Since $\int_{0}^{a}x^{n-\ell}\left(1-x\right)^{\ell-1}\,\text{d}x$ is of order $O(a^{n-\ell+1}n^{-\ell})$ and $\int_{a}^{1}x^{n-\ell-1}\left(1-x\right)^{\ell-1}\,\text{d}x$ is of order $O(n^{-\ell})$, we have that $M_{1}$ is of order $O(a)$. Similarly, since $\int_{a}^{1}x^{n-\ell-2}\left(1-x\right)^{\ell-1}\,\text{d}x$ is also of order $O(n^{-\ell})$, we have that $M_{2}$ is of order $O(a^{2})$.
Recalling that $\mathbb{E}[K_{n}(a,\ell)]-1=(n-\ell)M_{1}$, if we choose $a$ such that $an\to 0$ as $n\to\infty$, then
$\beta=\frac{(n-\ell)M_{1}}{(n-\ell)M_{1}+\ell}\to 0$
and $(n-\ell-1)\frac{M_{2}}{M_{1}}-(n-\ell-2)M_{1}\to 0$. Writing $(1-\beta)^{\ell}\geq 0$, Theorem 5 then gives us the upper bound
$d_{\text{TV}}(K_{n}(a,\ell)-1,Z)\leq\frac{(n-\ell)M_{1}+\ell}{\ell}\left(\beta+(1-\beta)\left[(n-\ell-1)\frac{M_{2}}{M_{1}}-(n-\ell-2)M_{1}\right]\right)\to 0$
as $n\to\infty$ for this choice of $a=a(n)$.